if you invest $5000 between two accounts paying 7% and 9% annual interest, respectively. If the total interest earned for the year was $430 how much was invested at each rest
Answer by addingup(3677)

.07x+.09(5000-x)= 430
.07x+450-.09x= 430
-.02x= -20
x= 1000 was invested at 7% and
5000-1000= 4000 was invested at 9%
Proof
1000*.07= 70
4000*.09= 360
_ _ _ _ _------
Total.....430 We have the correct answer.
J